OpenStudy (anonymous):

find θ if cotθ= -.577

OpenStudy (dumbcow):

\[\cot \theta = \frac{1}{\tan \theta} \rightarrow \tan \theta = \frac{1}{\cot \theta}\] \[\theta = \tan^{-1} (\frac{1}{-.577})\] use scientific calculator

OpenStudy (campbell_st):

well cot = 1/tan so then if \[\cot(\theta) = -0.577\] take the reciprocal of each and \[\tan(\theta) = \frac{1}{-0.577}\] and tan is negative in the 2nd quadrant and 4th quadrant so find \[\tan (\theta) = \frac{1}{0.577}\] then use 2nd quadrant \[ Angle = 180 - \theta\] 4th quadrant \[Angle = 360 - \theta\] hope it makes sense